find the odd no
5,6,19,41,105,______,446

Found 2 solutions by jorel1380, ikleyn:
Answer by jorel1380(3719)

5+1³=6
6+2³=14
14+3³=41
41+4³=105
105+5³=230
.
.
☺☺☺☺

Answer by ikleyn(52794)

.
find the odd no
5,6,19,41,105,______,446

~~~~~~~~~~~~~~~~~~


The solution by @jorell380 recalls me the old (and legendary) joke of physicists:

    "In this row the number 19 is an experimental error . . . "